# Compatible finite element interpolated neural networks

### v1

@@ Abstract.
We extend the finite element interpolated neural network (FEINN) framework from partial differential equations (PDEs) with weak solutions in $H^{1}$ to PDEs with weak solutions in $H(\textbf{curl})$ or $H(\textbf{div})$. To this end, we consider interpolation trial spaces that satisfy the de Rham Hilbert subcomplex, providing stable and structure-preserving neural network discretisations for a wide variety of PDEs. This approach, coined compatible FEINNs, has been used to accurately approximate the $H(\textbf{curl})$ inner product. We numerically observe that the trained network outperforms finite element solutions by several orders of magnitude for smooth analytical solutions. Furthermore, to showcase the versatility of the method, we demonstrate that compatible FEINNs achieve high accuracy in solving surface PDEs such as the Darcy equation on a sphere. Additionally, the framework can integrate adaptive mesh refinements to effectively solve problems with localised features. We use an adaptive training strategy to train the network on a sequence of progressively adapted meshes. Finally, we compare compatible FEINNs with the adjoint neural network method for solving inverse problems. We consider a one-loop algorithm that trains the neural networks for unknowns and missing parameters using a loss function that includes PDE residual and data misfit terms. The algorithm is applied to identify space-varying physical parameters for the $H(\textbf{curl})$ model problem from partial or noisy observations. We find that compatible FEINNs achieve accuracy and robustness comparable to, if not exceeding, the adjoint method in these scenarios.
@@ 1. Introduction
In this work, we integrate compatible FEs, i.e., spaces that form a discrete the Rham complex [Arnold2006], into the FEINN method proposed in [Badia2024]. This integration enables us to solve PDEs with weak solutions in $H(\textbf{curl})$ or $H(\textbf{div})$. We refer to this method as compatible FEINNs. The novelty of compatible FEINNs over the standard FEINN method in [Badia2024] lies in introducing curl/div-conforming trial FE spaces to interpolate NNs, ensuring the desired structure-preserving properties across element boundaries. The residual minimisation framework underlying FEINNs allows for some flexibility in selecting the trial and test FE spaces provided that an inf-sup compatibility condition is satisfied among them; see [Badia2024adaptive] for the numerical analysis of the method. As we showed in [Badia2024] and further confirmed in this paper for the problems at hand, a proper choice of these spaces can lead to significant improvements in the accuracy of the solution and the convergence of the optimiser. As an evidence on the soundness of this approach, we also showcase its applicability to solve PDEs posed over immersed manifolds. We interpolate NN vector-valued fields living in ambient space (and thus not necessarily tangent to the manifold) onto FE spaces made of functions constrained to be tangent to the manifold by construction. Strikingly, NNs trained on a coarse mesh and lower-order bases are able to provide several orders of magnitude higher accurate solutions to the problem when interpolated onto a FE space built out of finer meshes and higher order bases. Besides, this method, which resembles TraceFEM [Olshanskii2017] (as it seeks an approximation of the surface PDE on a finite-dimensional space of functions living on a higher dimensional space), does not need stabilisation terms in the loss function to enforce tangentiality, as it relies on a surface FE interpolation on the tangent space. A comprehensive set of numerical experiments is conducted to demonstrate the performance of compatible FEINNs in forward and inverse problems involving problems in $H(\textbf{curl})$ and $H(\textbf{div})$.
@@ 2.3.3. Linearised test spaces
Each hexahedral Nédélec element of order $k_{U}$ has a total of $3k_{U}(k_{U}+1)^{2}$ DoFs: $12k_{U}$ on the edges (with $k_{U}$ DoFs per edge), $6(2k_{U}^{2}-2k_{U})$ on the faces (with $2k_{U}^{2}-2k_{U}$ DoFs per face), and $3k_{U}(k_{U}-1)^{2}$ in the interior. After applying $k_{U}-1$ uniform refinements, the refined element has $k_{U}$ subedges on each of the original edges, $2k_{U}^{2}-2k_{U}$ subedges on each of the original faces, and $3k_{U}(k_{U}-1)^{2}$ interior subedges. The linearised Nédélec element also has $12k_{U}$ edge DoFs, $6(2k_{U}^{2}-2k_{U})$ face DoFs, and $3k_{U}(k_{U}-1)^{2}$ interior DoFs. Thus, the the DoFs per geometrical entity and dimensions of $U_{h}^{1}$ and ${V_{h}^{1}}$ are also equal for hexahedral Nédélec elements. ∎
@@ 2.9. Numerical analysis
where $U_{h}\subset U$ and ${V_{h}}\subset U$ are the discrete trial and test spaces, respectively. The discrete inf-sup condition, using $k_{U}=k_{V}$, is satisfied for the compatible FE spaces being used in this work, which is a direct consequence of the cochain projection between the continuous and discrete de Rham complexes. For the Petrov-Galerkin discretisation with linearised test spaces, i.e., $k_{U}>k_{V}=1$, we are not aware of any results in the literature. However, the experimental evaluation of condition numbers for the spaces being considered in this paper indicate that this inf-sup condition is satisfied for all the orders being considered in this work. Under these conditions, the following nonlinear Cea’s lemma holds, where we assume that the NN can emulate the FE space for simplicity (see [Badia2024adaptive] for a more general statement).
@@ 3. Numerical experiments
First, we use $h$-adaptive FEINNs to attack a problem with localised features, demonstrating that the trained NNs outperform FEM solutions without the need to use nested loops. In the inverse problem experiments, we focus on recovering the unknown physical parameters of Maxwell’s equations from partial or noisy observations, using the training strategy proposed in [Badia2024]. We compare the performance of compatible FEINNs with the adjoint NN method. With partial observations, the addition of an extra NN in the compatible FEINN method improves the accuracy of both the predicted parameters and recovered states compared to the adjoint NN method. When the observations are noisy, compatible FEINNs achieve comparable performance to the adjoint NN method without any regularisation terms in the loss function.
@@ 3.2. Forward problems
In addition to measuring the performance of the interpolated NNs, we are also interested in how well the NNs satisfy the Dirichlet boundary condition. We note that Dirichlet boundary condition is implicitly enforced to the NN via the FE residual minimisation. In each plot within this section, the label “FEINN” denotes results for the interpolated NNs, while the label “NN” indicates results for the NNs themselves. For reference, we display the FEM solution using Petrov-Galerkin discretisation in the plots, labelled as “FEM”.
@@ 3.2.1. Maxwell’s equation with a smooth solution
Since the variance for FEINN results with the same mesh size and $k_{U}$ is negligible, we display the mean values of these 10 runs. We observe that both the $L^{2}$ and $H(\textbf{curl})$ errors of FEINNs consistently fall on the corresponding FEM lines for both $k_{U}=1$ and $k_{U}=2$, indicating successful training and accurate FEINN solutions. For the NNs, the errors are always below the corresponding FEM lines, suggesting that the NNs not only accurately satisfies the Dirichlet boundary condition but also provide more accurate results than the FEM solutions. For example, in Fig. 1(a), with $h=0.01$ and $k_{U}=1$, the NN is more than 2 orders of magnitude more accurate than the FEM solution. Similar observations are found for the $H(\textbf{curl})$ error in Fig. 1(b). These findings extend the conclusion in [Badia2024] that NNs can outperform FEM solutions for the Poisson equation with smooth solutions to the Maxwell’s equations. For $k_{U}=2$, the difference between the NN and the FEM results is less pronounced than for $k_{U}=1$, but NNs still beat the FEM solutions. Besides, with a fine enough mesh, and $k_{U}=1$, it is remarkable to find that the convergence rate of the NNs matches the $k_{U}=2$ FEM rate. This suggests that the smoothness of the NNs can bring superconvergence in certain scenarios.
@@ 3.2.2. The effects of preconditioning on Maxwell’s equation
More interestingly, as shown in Fig. 2(b), with the $H(\textbf{curl})$ norm, the NN solution starts outperforming the FEMs solution after around 1,500 iterations and achieves nearly an order of magnitude lower $H(\textbf{curl})$ error by 6,000 iterations. The preconditioned PDE training serves as a post-processing step to improve FE solution accuracy (more than one order in the experiments) without increasing the mesh resolution or trial order. In summary, the initialisation step, while not required, can lower computational cost and provide faster error reduction. Moreover, the $\mathbf{B}_{\mathrm{lin}}$ preconditioner, combined with $L^{2}$ or $H(\textbf{curl})$ norms in the loss (17), greatly accelerates NN convergence and improves the accuracy of the trained NNs. These results suggest that these schemes can be effective for transient problems, where the NN will be properly initialised with the solution at the previous time step.
@@ 3.2.4. Adaptive training for Maxwell’s equation with localised features
We conclude the forward experiments section by employing the $h$-adaptive FEINN method proposed in [Badia2024adaptive] to tackle a forward Maxwell problem with localised features. We consider the following analytic solution defined in the unit square domain $\Omega=[0,1]^{2}$ with a pure Dirichlet boundary:
@@
Note that the x-component of $\mathbf{u}$ has a circular wave front centred at $(-0.05,-0.05)$ with a radius of 1.2, while the y-component features a sharp peak at $(0.5,0.5)$. These localised features are typically benchmarks for evaluating the performance of numerical methods equipped with adaptive mesh refinement.
We adopt the train, estimate, mark, and refine strategy proposed in [Badia2024adaptive] to adaptively refine the mesh and train FEINNs. Preconditioning is applied in the loss function to improve the training stability and to speed up the convergence. We use the $L^{2}$ norm in the preconditioned loss function (17) and, similar to [Badia2024adaptive] and the previous experiment, we utilise the preconditioner $\mathbf{B}_{\mathrm{lin}}$. We use different error indicators to guide mesh adaptation. Namely, the $H(\textbf{curl})$ error among the interpolated NN and the analytical solution, and the $L^{2}$ norm of the strong PDEs residual evaluated at the NN. These are labelled as “real” and “network”, respectively, in the figures below. We leverage forest-of-octrees meshes for adaptivity, which are typically non-conforming. To maintain curl-conforming property in the FE space, we incorporate additional linear multi-point constraints for slave DoFs. For a more detailed discussion on $h$-adaptive FEINN, we refer to [Badia2024adaptive].
@@
Figure 6. Error convergence of NN solutions with respect to refinement steps for the forward Maxwell problem with localised features, using different error indicators. The solid line denotes the median, and the band represents the range from the minimum to the 90th percentile across 20 independent runs. The red dashed line illustrates the $h$-adaptive FEM solution error using the real FEM $H(\textbf{curl})$ error as the refinement indicator.
Although the trained NNs start with higher errors, they begin to outperform the $h$-adaptive FEM solution after the first two refinement steps. As shown in both Fig. 6(a) and Fig. 6(b), the NN errors consistently fall below the FEM error lines after the second refinement. This observation agrees with the findings in [Badia2024adaptive], where NNs can beat $h$-adaptive FEM for smooth analytic solutions. However, in [Badia2024adaptive], NN solutions are noted to be superior mainly in $H^{1}$-error for Poisson equations, but our Maxwell experiments show lower errors in both $L^{2}$ and $H(\textbf{curl})$ norms. Furthermore, the network error indicator effectively guides mesh refinements, as reflected by the closely aligned convergence curves across different indicators. A similar finding was also reported in [Badia2024adaptive] for the Poisson problems.

### v2

@@ Abstract.
We extend the finite element interpolated neural network (FEINN) framework from partial differential equations (PDEs) with weak solutions in $H^{1}$ to PDEs with weak solutions in $H(\textbf{curl})$ or $H(\textbf{div})$. To this end, we consider interpolation trial spaces that satisfy the de Rham Hilbert subcomplex, providing stable and structure-preserving neural network discretisations for a wide variety of PDEs. This approach, coined compatible FEINNs, has been used to accurately approximate the $H(\textbf{curl})$ inner product. We numerically observe that the trained network outperforms finite element solutions by several orders of magnitude for smooth analytical solutions. Furthermore, to showcase the versatility of the method, we demonstrate that compatible FEINNs achieve high accuracy in solving surface PDEs such as the Darcy equation on a sphere. Additionally, the framework can integrate adaptive mesh refinements to effectively solve problems with localised features. We use an adaptive training strategy to train the network on a sequence of progressively adapted meshes. Finally, we compare compatible FEINNs with the adjoint neural network method for solving inverse problems. We consider a one-loop algorithm that trains the neural networks for unknowns and missing parameters using a loss function that includes PDE residual and data misfit terms. The algorithm is applied to identify space-varying physical parameters for the $H(\textbf{curl})$ model problem from partial, noisy, or boundary observations. We find that compatible FEINNs achieve accuracy and robustness comparable to, if not exceeding, the adjoint method in these scenarios.
@@ 1. Introduction
In this work, we integrate compatible FEs, i.e., spaces that form a discrete the Rham complex [Arnold2006], into the FEINN method proposed in [Badia2024]. This integration enables us to solve PDEs with weak solutions in $H(\textbf{curl})$ or $H(\textbf{div})$. We refer to this method as compatible FEINNs. The novelty of compatible FEINNs over the standard FEINN method in [Badia2024] lies in introducing curl/div-conforming trial FE spaces to interpolate NNs, ensuring the desired structure-preserving properties across element boundaries. The residual minimisation framework underlying FEINNs allows for some flexibility in selecting the trial and test FE spaces provided that an inf-sup compatibility condition is satisfied among them; see [Badia2024adaptive] for the numerical analysis of the method. As we showed in [Badia2024] and further confirmed in this paper for the problems at hand, a proper choice of these spaces can lead to significant improvements in the accuracy of the solution and the convergence of the optimiser. As an evidence on the soundness of this approach, we also showcase its applicability to solve PDEs posed over immersed manifolds. We interpolate NN vector-valued fields living in ambient space (and thus not necessarily tangent to the manifold) onto FE spaces made of functions constrained to be tangent to the manifold by construction. Strikingly, NNs trained on a coarse mesh and lower-order bases are able to provide several orders of magnitude higher accurate solutions to the problem when interpolated onto a FE space built out of finer meshes and higher order bases. Besides, this method, which resembles TraceFEM [Olshanskii2017] (as it seeks an approximation of the surface PDE on a finite-dimensional space of functions living on a higher dimensional space), does not need stabilisation terms in the loss function to enforce tangentiality, as it relies on a surface FE interpolation on the tangent space. A comprehensive set of numerical experiments is conducted to demonstrate the performance of compatible FEINNs in forward and inverse problems involving PDEs in $H(\textbf{curl})$ and $H(\textbf{div})$.
@@ 2.3.3. Linearised test spaces
Each hexahedral Nédélec element of order $k_{U}$ has a total of $3k_{U}(k_{U}+1)^{2}$ DoFs: $12k_{U}$ on the edges (with $k_{U}$ DoFs per edge), $6(2k_{U}^{2}-2k_{U})$ on the faces (with $2k_{U}^{2}-2k_{U}$ DoFs per face), and $3k_{U}(k_{U}-1)^{2}$ in the interior. After applying $k_{U}-1$ uniform refinements, the refined element has $k_{U}$ subedges on each of the original edges, $2k_{U}^{2}-2k_{U}$ subedges on each of the original faces, and $3k_{U}(k_{U}-1)^{2}$ interior subedges. The linearised Nédélec element also has $12k_{U}$ edge DoFs, $6(2k_{U}^{2}-2k_{U})$ face DoFs, and $3k_{U}(k_{U}-1)^{2}$ interior DoFs. Thus, the DoFs per geometrical entity and dimensions of $U_{h}^{1}$ and ${V_{h}^{1}}$ are also equal for hexahedral Nédélec elements. ∎
@@ 2.7. Loss functions
After interpolation, $\pi^{1}_{h,0}(\mathbf{u}_{\mathcal{N}})\in U^{1}_{h,0}$ vanishes on $\Gamma_{D}$, ensuring that the FEINN solution $\pi^{1}_{h,0}(\mathbf{u}_{\mathcal{N}})+\bar{\mathbf{u}}_{h}$ strongly satisfies the Dirichlet boundary conditions. In FEM, $\bar{\mathbf{u}}_{h}$ is a standard component and can be computed efficiently. Our method provides a simple and effective way to strongly enforce these conditions, avoiding the need for penalty terms in the loss function [Raissi2019], or the use of distance functions [Sukumar2022, Berrone2022]. Additionally, Gaussian quadrature rules allow efficient and exact computation of the integrals in (17). Spatial derivatives are now calculated using the polynomial FE basis functions, which are more efficient than hybrid automatic differentiation (mixed derivatives with respect to parameters and inputs) widely used in standard PINN methods. Furthermore, the use of a residual dual function (i.e., the inclusion of preconditioners) further ensures well-posedness of the problem and accelerates training convergence.
@@ 2.9. Numerical analysis
where $U_{h}\subset U$ and ${V_{h}}\subset U$ are the discrete trial and test spaces, respectively. The discrete inf-sup condition, using $k_{U}=k_{V}$, is satisfied for the compatible FE spaces being used in this work, which is a direct consequence of the cochain projection between the continuous and discrete de Rham complexes. Using other FE spaces, such as Lagrange elements, fails the discrete inf-sup condition, making the discrete problem ill-posed. In such cases, regularisation is required to ensure a well-posed FE problem. See [Costabel2002, Badia2012-xf, Badia2009-yj] for more details. For the Petrov-Galerkin discretisation with linearised test spaces, i.e., $k_{U}>k_{V}=1$, we are not aware of any results in the literature. However, the experimental evaluation of condition numbers for the spaces being considered in this paper indicate that this inf-sup condition is satisfied for all the orders being considered in this work. Under these conditions, the following nonlinear Cea’s lemma holds, where we assume that the NN can emulate the FE space for simplicity (see [Badia2024adaptive] for a more general statement).
@@ 3. Numerical experiments
First, we use $h$-adaptive FEINNs to attack problems with localised features, demonstrating that the trained NNs outperform FEM solutions for sharp-feature analytic solutions and match FEM accuracy for singular solutions, all without nested loops. In the inverse problem experiments, we focus on recovering the unknown physical parameters of Maxwell’s equations from partial or noisy observations, using the training strategy proposed in [Badia2024]. We compare the performance of compatible FEINNs with the adjoint NN method. With partial observations, the addition of an extra NN in the compatible FEINN method improves the accuracy of both the predicted parameters and recovered states compared to the adjoint NN method. When the observations are noisy, compatible FEINNs achieve comparable performance to the adjoint NN method without any regularisation terms in the loss function.
We omit the comparison of the FEINN method with other NN-based approaches, such as PINNs and VPINNs. This comparison is already studied in detail in [Berrone2022, Badia2024]. These studies find that the interpolated NN approach performs similar to or better than PINNs or VPINNs for a given number of NN evaluations. Additionally, the computational cost of interpolated NNs is lower than that of standard PINN methods, as spatial derivatives of NNs are handled using polynomial interpolation. Furthermore, as demonstrated in [Badia2024, Badia2024adaptive], FEINN convergence can be significantly improved by using preconditioned losses.
@@ 3.2. Forward problems
In addition to measuring the performance of the interpolated NNs, we are also interested in how well the NNs satisfy the Dirichlet boundary condition. We note that Dirichlet boundary condition is implicitly enforced to the NN via the FE residual minimisation. After the training process, in which the interpolated NN is used to compute the loss and gradients, we obtain a NN solution $\mathbf{u}_{\mathcal{N}}$. With this solution, we can evaluate the error of the interpolated NN, i.e., $\pi_{h,0}(\mathbf{u}_{\mathcal{N}})+\bar{\mathbf{u}}_{h}-\mathbf{u}$, or the NN itself, i.e., $\mathbf{u}_{\mathcal{N}}-\mathbf{u}$, where $\mathbf{u}$ is the true solution. In each plot within this section, the label “FEINN” denotes results for the interpolation of the NN solutions, while the label “NN” indicates results for the NNs themselves (using very accurate quadratures to approximate integrals of NNs). For reference, we display the FEM solution using Petrov-Galerkin discretisation in the plots, labelled as “FEM”.
@@ 3.2.1. Maxwell’s equation with a smooth solution
Since the variance for FEINN results with the same mesh size and $k_{U}$ is negligible, we display the mean values of these 10 runs. We observe that both the $L^{2}$ and $H(\textbf{curl})$ errors of FEINNs consistently fall on the corresponding FEM lines for both $k_{U}=1$ and $k_{U}=2$, indicating successful training and accurate FEINN solutions. For the NNs, the errors are always below the corresponding FEM lines, suggesting that the NNs not only accurately satisfies the Dirichlet boundary condition but also provide more accurate results than the FEM solutions. For example, in Fig. 1(a), with $h=0.01$ and $k_{U}=1$, the NN is more than 2 orders of magnitude more accurate than the FEM solution Similar observations are found for the $H(\textbf{curl})$ error in Fig. 1(b). These findings extend the conclusion in [Badia2024] that NNs can outperform FEM solutions for the Poisson equation with smooth solutions to the Maxwell’s equations. For $k_{U}=2$, the difference between the NN and the FEM results is less pronounced than for $k_{U}=1$, but NNs still beat the FEM solutions. Besides, with a fine enough mesh, and $k_{U}=1$, it is remarkable to find that the convergence rate of the NNs matches the $k_{U}=2$ FEM rate. This suggests that NNs can bring superconvergence in certain scenarios. The improved accuracy of the NNs compared to the FEM solutions is likely attributed to the smoothness of the NN solutions: the $\mathcal{C}^{\infty}$ NN solution (with the tanh activation function) outperforms the $\mathcal{C}^{0}$ FEM solution when approximating a smooth true state.
@@ 3.2.2. The effects of preconditioning on Maxwell’s equation
More interestingly, as shown in Fig. 2(b), with the $H(\textbf{curl})$ norm, the NN solution starts outperforming the FEMs solution after around 1,500 iterations and achieves nearly an order of magnitude lower $H(\textbf{curl})$ error by 6,000 iterations. The preconditioned PDE training serves as a post-processing step to improve FE solution accuracy (more than one order of magnitude in the experiments) without increasing the mesh resolution or trial order. In summary, the initialisation step, while not required, can lower computational cost and provide faster error reduction. Moreover, the $\mathbf{B}_{\mathrm{lin}}$ preconditioner, combined with $L^{2}$ or $H(\textbf{curl})$ norms in the loss (17), greatly accelerates NN convergence and improves the accuracy of the trained NNs. These results suggest that these schemes can be effective for transient problems, where the NN will be properly initialised with the solution at the previous time step.
@@ 3.2.4. Adaptive training for Maxwell’s equation with localised features
We conclude the forward experiments section by employing the $h$-adaptive FEINN method proposed in [Badia2024adaptive] to tackle forward Maxwell problems with localised features. We consider both sharp features and singularities in the solution. We first focus on the following analytic solution defined in the unit square domain $\Omega=[0,1]^{2}$ with a pure Dirichlet boundary:
@@
Note that the x-component of $\mathbf{u}$ has a circular wave front centred at $(-0.05,-0.05)$ with a radius of 1.2, while the y-component features a sharp peak at $(0.5,0.5)$. These sharp features are typically benchmarks for evaluating the performance of numerical methods equipped with adaptive mesh refinement.
We adopt the train, estimate, mark, and refine strategy proposed in [Badia2024adaptive] to adaptively refine the mesh and train FEINNs. Preconditioning is applied in the loss function to improve the training stability and to speed up the convergence. We use the $L^{2}$ norm in the preconditioned loss function (17) and, similar to [Badia2024adaptive] and the previous experiment, we utilise the preconditioner $\mathbf{B}_{\mathrm{lin}}$. We use different error indicators to guide mesh adaptation. Namely, the $L^{2}$ error among the interpolated NN and the analytical solution, the $L^{2}$ norm of the difference between the NN and its FE interpolation, and the $L^{2}$ norm of the strong PDEs residual evaluated at the NN. These are labelled as “real”, “integration”, and “network”, respectively, in the figures below. We leverage forest-of-octrees meshes for adaptivity, which are typically non-conforming. To maintain curl-conforming property in the FE space, we incorporate additional linear multi-point constraints for slave DoFs. For a more detailed discussion on $h$-adaptive FEINN, we refer to [Badia2024adaptive].
@@
Figure 6. Error convergence of NN solutions with respect to refinement steps for the forward Maxwell problem with less smooth solution, using different error indicators. The solid line denotes the median, and the band represents the range from the minimum to the 90th percentile across 20 independent runs. The red dashed line illustrates the $h$-adaptive FEM solution error using the real FEM $H(\textbf{curl})$ error as the refinement indicator.
Although the trained NNs start with higher errors, they begin to outperform the $h$-adaptive FEM solution after the first two refinement steps. As shown in both Fig. 6(a) and Fig. 6(b), the NN errors consistently fall below the FEM error lines after the second refinement. This observation agrees with the findings in [Badia2024adaptive], where NNs can beat $h$-adaptive FEM for smooth analytic solutions. However, in [Badia2024adaptive], NN solutions are noted to be superior mainly in $H^{1}$-error for Poisson equations, but our Maxwell experiments show lower errors in both $L^{2}$ and $H(\textbf{curl})$ norms. Furthermore, the network and integration error indicators effectively guides mesh refinements, as reflected by the closely aligned convergence curves across different indicators. A similar finding was also reported in [Badia2024adaptive] for the Poisson problems.
Next, we increase the difficulty of the problem by introducing singular feature in the solution. The analytic solution is now defined in the domain $\Omega=[0,1]^{2}$ as follows:
$\mathbf{u}(x,y)=\boldsymbol{\nabla}(r^{\frac{2}{3}}\sin(\frac{2\theta}{3})),\quad r=\sqrt{x^{2}+y^{2}},\quad\theta=\arctan(\frac{y}{x}).$
This solution exhibits a singularity at the origin and serves as a popular benchmark for testing adaptive methods (see, e.g., [Olm2019]). Most of the settings remain the same as the previous experiment, except for the following adjustments: the initial mesh comprises $20\times 20$ quadrilateral elements, the $H(\textbf{curl})$ norm is used in the preconditioned loss function, and the training iterations are in total 5,000 with $[500,750,1000,1250,1500]$ iterations for each step.
Fig. 7 displays the convergence of the NN errors across refinement step from 20 independent runs. The NN errors are computed by interpolating the trained NNs onto a second-order FE space on a finer uniform mesh. Despite the low regularity, the NN solutions consistently match the adaptive FEM solutions at each refinement step for both $L^{2}$ and $H(\textbf{curl})$ errors. The figure also highlights that the proposed network and integration error indicators effectively guide mesh refinements for this singular problem, achieving performance comparable to the real error that is barely available in practice. Overall, Fig. 6 and Fig. 7 illustrate the effectiveness of the $h$-adaptive FEINN method in solving $H(\textbf{curl})$ problems with sharp features and singularities.
((a))
((b))
Figure 7. Error convergence of NN solutions with respect to refinement steps for the forward Maxwell problem with singularity, using different error indicators. See the caption of Fig. 6 for details about the information displayed in this figure.